# Front stability of infinitely steep travelling waves in population biology

### 21 Dec 2023 v1

@@ 3.3 Travelling wavefront stability: Numerical explorations
Profiles in Figure 5(a)–(c) show three travelling wave solutions for $\kappa=0.1$ with $m=0.5,1.0,2.0$, respectively, where a small transverse perturbation is added. We solve Equations (2a)–(2d) with the level–set method with the profiles in Figure 5(a)–(c) as the initial condition. These solutions evolve to the profiles shown in Figure 5(d)–(f) at $t=50$ where we see that the amplitude of the perturbation has decayed so that the moving boundary is almost perfectly vertical. Additional profiles in Figure 5(g)–(i) show the time evolution of $\partial\Omega(t)$. Comparing the evolution of the shape of the moving boundary indicates that the amplitude of the initial perturbation decays with time indicating that these invading travelling wave profiles are stable when a small transverse perturbation is added.
Figure 5: Numerical exploration of front stability with $\kappa=0.1$. Profiles in (a)–(c) show three perturbed travelling wave profiles for $m=0.5,1.0,2.0$, respectively. Each perturbed travelling wave profile is computed by first solving the $\mathcal{O}(1)$ boundary value problem to give the shape of the unperturbed travelling wave that is then shifted so that $\partial\Omega(0)$ coincides with the vertical line at $x=5$ before a perturbation with wavenumber $q=4\pi/5$ and amplitude $\varepsilon=0.1$ is added to give the sinusoidal profiles in (a)–(c) at $t=0$. Results in (d)–(f) show the solution of Equations (2a)–(2d) at $t=50$ where, in each case, we can see that each front moves in the positive $x$–direction and amplitude of the transverse perturbation clearly decays. Plots in (g)–(i) show the evolution of $\partial\Omega(t)$ at $t=0,2,4,6,8$ with the arrow showing the increasing value of $t$. Comparing these plots of $\partial\Omega(t)$ suggests that these travelling waves with $\kappa=0.1$ are stable as the perturbation amplitude appears to decay with time. All numerical calculations are performed on a $10\times 10$ domain uniformly discretised with a $201\times 201$ mesh.
@@
Figure 6: Numerical exploration of front stability with $\kappa=-0.1$. Profiles in (a)–(c) show three perturbed travelling wave profiles for $m=0.5,1.0,2.0$, respectively. Each perturbed travelling wave profile is computed by first solving the $\mathcal{O}(1)$ boundary value problem to give the shape of the unperturbed travelling wave that is then shifted so that $\partial\Omega(0)$ coincides with the vertical line at $x=5$ before a perturbation with wavenumber $q=4\pi/5$ and amplitude $\varepsilon=0.1$ is added to give the sinusoidal profiles in (a)–(c) at $t=0$. Results in (d)–(f) show the solution of Equations (2a)–(2d) at $t=50$ where, in each case, we can see that each front moves in the positive $x$–direction and amplitude of the transverse perturbation clearly decays. Plots in (g)–(i) show the evolution of $\partial\Omega(t)$ at $t=0,2,4,6,8$ with the arrow showing the increasing value of $t$. Comparing these plots of $\partial\Omega(t)$ suggests that these travelling waves with $\kappa=0.1$ are stable as the perturbation amplitude appears to grow with time. All numerical calculations are performed on a $10\times 10$ domain uniformly discretised with a $201\times 201$ mesh.
@@ 3.4 Travelling wavefront stability: Linear stability analysis
were, $q$ is the wave number and $\varepsilon\ll 1$ is the amplitude of the small perturbation Muller2002 ; Tam2023 . Here $\omega$ is an unknown constant that we will use to describe the short time decay $\omega<0$ or growth $\omega>0$ of the transverse perturbation. Our aim now is to assume that we have $\varepsilon\ll 1$ and use this small parameter to derive a perturbation solution to explore the dispersion relationship $\omega(q)$ for a range of travelling wave solutions characterised by different choices of $c$ and $m$ which, as we showed in Section 3.2 implicitly defines the associated travelling wave speed. This approach involves numerically solving a series of boundary value problems with one independent variable for the real component of the growth rate $\omega$.
@@
To compute these quantities we use linear interpolation to find positions $x_{j}$ such that $\phi(x,y_{j},t_{k})=0$ for the $j$th row in the discretisation of $\mathcal{D}$ at time $t=t_{k}$. The maximum and minimum values in the set ${x_{j}}$ determine $X_{\textrm{max}}(t)$ and $X_{\textrm{min}}(t)$, respectively. To estimate $\omega$ we use Equation (16) and assume that the perturbation amplitude grows or decays exponentially fast over a sufficiently short time interval. We use the Polynomials.jl package in Julia to perform a least–squares fit of $\log(A)=\omega_{\textrm{n}}t+C$, where $C$ is a constant, and $\omega_{\textrm{n}}$ is the numerical estimate of the growth rate. For all results presented we consider a relatively short interval $t\in[1,2]$ with 21 equally–spaced time points within that interval to compute $\omega_{\textrm{n}}$. For each problem considered we generate a plot of $\log(A)$ as a function of $t$ which allows us to check that $A(t)$ grows or decays exponentially with time, as indicated by our data falling on a straight line over the time interval considered.
@@
Results in Figure 67 allow us to briefly return to Section 1 where, in Figure 1(e)–(h) where we hypothesized four potential outcomes in terms of invading, receding, linearly stable and linearly unstable travelling waves. Our results in Figure 78 indicate that invading travelling waves with $\kappa=0.1$ and $m=1$ are linearly stable while receding travelling waves with $\kappa=-0.1$ and $m=1$ are linearly unstable. We have found that these broad hold for other choices of $m$ and $\kappa$, namely that setting $\kappa>0$ leads to invading travelling waves that are linearly stable and setting $\kappa<0$ leads to receding travelling waves that are linearly unstable. Our open-source Julia codes provided on Github can be used to explore these other parameter choices. In summary, of the four potential outcomes previously hypothesized in Figure 1(e)–(h) we have only found evidence to support the scenario depicted in Figure 1(e) for $\kappa>0$ and Figure 1(h) for $\kappa<0$.
@@ 4 Conclusion and future work
In this work we re–visited the recently–established Porous–Fisher–Stefan (PFS) model of biological invasion, and focused our attention on properties of the travelling wave solutions of that model. The first presentation of the PFS model focused on working with a degenerate nonlinear diffusion term with flux $\mathcal{J}=-u\partial u/\partial x$ on $-\infty<x<L(t)$ with $\textrm{d}L(t)/\textrm{d}t=-\kappa u\partial u/\partial x$ at $x=L(t)$, and the existence of travelling waves describing biological invasion with $\kappa>0$ were established using phase plane and perturbation methods Fadai2020 . Here we extend this modelling framework by dealing with a 2D model that incorporates a generalised degenerate diffusion term with flux $-u^{m}\nabla u$, and we consider both invading fronts with $\kappa>0$, and receding fronts with $\kappa<0$. Solving the 2D model in an appropriate domain provides strong numerical evidence that the model supports long–time travelling wave solutions, and working in the phase plane allows us to examine properties of the travelling wave solutions over a range of model parameters in the PDE model, and to determine how the long–time travelling wave speed $c$ depends upon the choice of parameters in the PDE model.
@@
Preliminary explorations of the time–dependent PDE solutions indicate that invading travelling waves with $\kappa>0$ and $c>0$ are stable when a small transverse perturbation with wavenumber $q$ is introduced at the moving boundary, whereas receding travelling waves with $\kappa<0$ and $c<0$ are unstable. These informal numerical observations are corroborated with a linear stability analysis indicating that the short–time amplitude of the transverse perturbation increases when $\kappa<0$, whereas it decreases with $\kappa>0$. In general we find reasonably good agreement between the dispersion relationship obtained from the linear stability analysis and from the 2D level–set method, especially for small $q$. For larger $q$ it is straightforward to carry out the linear stability analysis because this involves solving two boundary value problems with just one independent variable that we discretise efficiently using a non-uniform mesh. In contrast, solving the 2D time–dependent PDE problem for larger $q$ requires and increasingly refined spatial mesh as $q$ increases, which makes the time–dependent PDE solutions impractical.
@@
As illustrated in Figures 78 setting $\kappa<0$ leads to linearly unstable travelling wave solutions. In the heat transfer literature, linearly unstable front propagation is often studied by regularisation by including additional mechanisms in the PDE model that can stabilise perturbation modes with large wavenumbers Chadam1983 ; Yang2002 . A standard approach in heat and mass transfer for Stefan–type moving boundary problems is to include terms describing surface tension effects Chadam1983 . Introducing surface tension involves replacing the usual 1D interface condition $u(L(t),t)=0$ with $u(L(t),t)=\gamma\mathcal{K}$, where $\gamma>0$ is the surface tension coefficient, and $\mathcal{K}$ is the signed curvature of the moving front Chadam1983 . This approach is reasonable in heat transfer and fluid mechanics where the dependent variable represents temperature and pressure, respectively. This means that the boundary condition $u(L(t),t)=0$ corresponds to a zero scaled temperature or zero scaled reference pressure, and it is physically reasonable to consider a negative scaled temperature or pressure at the 2D moving front with negative curvature. Working with the PFS model is very difference since here $u$ represents a scaled population density which means that it is physically unreasonable to set $u(L(t),t)=\gamma\mathcal{K}$ which would be negative at the 2D moving front with negative curvature. Therefore, in this work we have not pursued any regularisation of the linearly unstable travelling wave fronts as the classical approach is not suitable for this application.
@@
There are many opportunities for future work motivated by the problems that we have considered here. For example, one question for future consideration is to explore different options for biologically meaningful regularisation terms that could be applied both to linearly unstable front propagation for the PFS model, as well as other moving boundary problems that describe population biology phenomena. Other technical challenges could also be considered. For example, our phase plane analysis in Section 3.2 led to expressions for an exact solution for $c=0$ and $m=1$, which we used to construct perturbation solutions for $|c|\ll 1$. These solutions are insightful because they can be used to provide insight into the relationship between $\kappa$ and $c$ for $m=1$. Another option would be to use these solutions to construct perturbation solutions for $|m-1|\ll 1$ which could be used to provide insight into the relationship between $\kappa$ and $c$ for $m\neq 1$. We leave both these extensions for future consideration.
@@
We solve Equations (19a)–(19d) by re–writing the boundary value problem in terms of $\theta_{0}=u_{0}^{(1+m)}$,
$\theta_{0}^{m/(1+m)}\left(\frac{\mathrm{d}^{2}\theta_{0}}{\mathrm{d}\xi^{2}}\right)+c\;\left(\frac{\mathrm{d}\theta_{0}}{\mathrm{d}\xi}\right)+(1+m)\;\theta_{0}\left(1-\theta_{0}^{1/(1+m)}\right)=0,\;\;\text{on}\;\;-\infty<\xi<0,$ (23a)
$\lim_{\xi\to-\infty}\theta_{0}(\xi)=1,$ (23b)
$\theta_{0}=0,\;\;\text{at}\;\;\xi=0,$ (23c)
$\frac{\mathrm{d}\theta_{0}}{\mathrm{d}\xi}$ $=-\frac{(1+m)\;c}{\kappa}\;\;\text{at}\;\;\xi=0.$ (23d)
@@
$\theta_{0}^{m/(1+m)}\left(\frac{\mathrm{d}^{2}\theta_{0}}{\mathrm{d}\xi^{2}}\right)+c\;\left(\frac{\mathrm{d}\theta_{0}}{\mathrm{d}\xi}\right)+(1+m)\;\theta_{0}\left(1-\theta_{0}^{1/(1+m)}\right)=0,\;\;\text{on}\;\;-\xi_{\textrm{max}}<\xi<0,$ (24a)
$\theta_{0}(-\xi_{\textrm{max}})=1,$ (24b)
$\theta_{0}=0,\;\;\text{at}\;\;\xi=0,$ (24c)
$\frac{\mathrm{d}\theta_{0}}{\mathrm{d}\xi}$ $=-\frac{(1+m)\;c}{\kappa}\;\;\text{at}\;\;\xi=0,$ (24d)
@@
In general $\theta_{0}$ varies rapidly with $\xi$ near $\xi=0$, whereas $\theta_{0}$ approaches unity as $\xi\to-\xi_{\textrm{max}}$. We take advantage of these properties by employing a non-uniform mesh where the mesh resolution is fine near $\xi=0$ and becomes increasingly coarse near $\xi=-\xi_{\textrm{max}}$. One way of achieving this is to use a geometrically spaced mesh McCue2022 where the mesh spacing varies from $\delta\xi_{\textrm{min}}$ at $\xi=0$ to $\delta\xi_{\textrm{max}}$ at $\xi=-\xi_{\textrm{max}}=-20$ using $N$ geometrically–spaced mesh points. At the $i$th internal mesh points we have $\delta\xi_{i}=(1+\epsilon)\delta\xi_{i+1}$. For $\delta\xi_{\textrm{min}}=1\times 10^{-5}$ with $N=301$ mesh points we obtain a geometric expansion factor of $(1+\epsilon)=1.0381894$ and $\delta\xi_{\textrm{max}}=0.7357$. Numerical experimentation confirms that this nonuniform mesh spacing leads to grid–independent results for the problems that we consider.
@@
At the $i$th internal mesh point we define $h^{+}_{i}=\xi_{i+1}-\xi_{i}$ and $h^{-}_{i}=\xi_{i}-\xi_{i-1}$ and discretise Equations(24a)–(24c) on the non-uniform mesh leading to a system of $N$ nonlinear algebraic equations $\boldsymbol{F}(\theta_{0}^{(i)})=\boldsymbol{0}$ which we solve using Newton-Raphson iteration. Elements of the vector $\boldsymbol{F}$ are given by,
$F_{0}$ $=\theta_{0}^{(0)}-1,$ (25a)
$F_{i}$ $={\theta_{0}^{(i)}}^{m/(1+m)}\left(A\theta_{0}^{(i+1)}+B\theta_{0}^{(i)}+C\theta_{0}^{(i-1)}\right)+c\frac{\theta_{0}^{(i+1)}-\theta_{0}^{(i-1)}}{h_{i}^{+}+h_{i}^{-}}+(1+m)\theta_{0}^{(i)}\left(1-{\theta_{0}^{(i)}}^{1/(1+m)}\right),$ (25b)
$F_{N}$ $=\theta_{0}^{(N)}\;,$ (25c)
@@
where the superscript $i=2,\dots,N-1$ refers to the $i$th internal mesh point, and the coefficients $A={2}/({h^{+}_{i}(h_{i}^{+}+h_{i}^{-})})$, $B={-2}/({h_{i}^{+}h_{i}^{-}})$, and $C={2}/({h^{-}_{i}(h_{i}^{+}+h_{i}^{-})})$ arise by applying standard finite difference stencils on the nonuniform mesh McCue2022 . An initial estimate of $\boldsymbol{\theta}_{0}^{(0)}$ can be updated using Newton-Raphson iteration as follows,
$\boldsymbol{\theta}_{0}^{(n+1)}=\boldsymbol{\theta}_{0}^{(n)}-J^{-1}\left(\boldsymbol{\theta}_{0}^{(n)}\right)\boldsymbol{F}\left(\boldsymbol{\theta}_{0}^{(n)}\right),$ (26)
@@
where the superscript $n$ gives the iteration count and $J_{ij}=\partial_{j}F_{i}$ are the elements of the Jacobian matrix for $\boldsymbol{F}(\theta_{0}^{(i)})$. Iterations are terminated when $\lVert\boldsymbol{\theta}_{0}^{(n+1)}-\boldsymbol{\theta}_{0}^{(n)}\rVert<1\times 10^{-6}$. In this work we take $\boldsymbol{\theta}_{0}^{(0)}$ to be a piecewise linear approximation of the shape of the travelling wave profile.
@@
To estimate the travelling wave speed $c$ we use a shooting method that commences with an initial estimate of $c^{(0)}$ to solve the leading order boundary value problem for $\theta_{0}(\xi)$. For our estimates of $c^{(0)}$ and $\theta_{0}(\xi)$ we discretise the moving boundary condition, Equation (23d), on our non-uniform mesh to give
$f^{(k)}=c^{(k)}+\frac{\kappa}{m+1}\left(P\theta_{0}^{({N-2},k)}-Q\theta_{0}^{({N-1},k)}+R\theta_{0}^{({N},k)}\right),$ (27)
@@
where $P=h^{+}_{N-1}/({h^{-}_{N-1}(h^{+}_{N-1}+h^{-}_{N-1})})$, $Q=(h^{+}_{N-1}+h^{-}_{N-1})/(h^{+}_{N-1}h^{-}_{N-1})$,, and $R=(2h^{+}_{N-1}+h^{-}_{N-1})/({h^{+}_{N-1}(h^{+}_{N-1}+h^{-}_{N-1})})$ is the estimate of the wave speed after $k$ iterations and while $\theta_{0}^{(i,k)}$ denotes our estimate of $\theta_{0}$ on the $i$th mesh point after $k$ iterations of the shooting method. To satisfy Equation (23d) we update our estimate of $c$ so that $f=0$, again using Newton-Raphson iteration,

### 06 Jan 2024 v2

@@ 3.3 Travelling wavefront stability: Numerical explorations
Profiles in Figure 5(a)–(c) show three travelling wave solutions for $\kappa=0.1$ with $m=0.5,1.0,2.0$, respectively, where a small transverse perturbation is added in each case. We solve Equations (2a)–(2d) with the level–set method with the profiles in Figure 5(a)–(c) as the initial condition. These solutions evolve to the profiles shown in Figure 5(d)–(f) at $t=50$ where we see that the amplitude of the perturbation decays so that the moving boundary is almost perfectly vertical. Additional profiles in Figure 5(g)–(i) show the time evolution of $\partial\Omega(t)$. Comparing the evolution of the shape of the moving boundary indicates that the amplitude of the initial perturbation decays with time indicating that these invading travelling wave profiles are stable when a small transverse perturbation is added.
Figure 5: Numerical exploration of front stability with $\kappa=0.1$. Profiles in (a)–(c) show three perturbed travelling wave profiles for $m=0.5,1.0,2.0$, respectively. Each perturbed travelling wave profile is computed by first solving the $\mathcal{O}(1)$ boundary value problem to give the shape of the unperturbed travelling wave that is then shifted so that $\partial\Omega(0)$ coincides with the vertical line at $x=5$ before a perturbation with wavenumber $q=4\pi/5$ and amplitude $\varepsilon=0.1$ is added to give the sinusoidal profiles in (a)–(c) at $t=0$. Results in (d)–(f) show the solution of Equations (2a)–(2d) at $t=50$ where, in each case, we can see that each front moves in the positive $x$–direction and amplitude of the transverse perturbation clearly decays. Plots in (g)–(i) show the evolution of $\partial\Omega(t)$ at $t=0,2,4,6,8$ with the arrow showing the increasing value of $t$. Note that the plots in (g)–(i) focus on the interval $4\leq x\leq 6$ so that the details of the interface evolution are clear. Comparing these plots of $\partial\Omega(t)$ suggests that these travelling waves with $\kappa=0.1$ are stable as the perturbation amplitude appears to decay with time. All numerical calculations are performed on a $10\times 10$ domain uniformly discretised with a $201\times 201$ mesh.
@@
Figure 6: Numerical exploration of front stability with $\kappa=-0.1$. Profiles in (a)–(c) show three perturbed travelling wave profiles for $m=0.5,1.0,2.0$, respectively. Each perturbed travelling wave profile is computed by first solving the $\mathcal{O}(1)$ boundary value problem to give the shape of the unperturbed travelling wave that is then shifted so that $\partial\Omega(0)$ coincides with the vertical line at $x=5$ before a perturbation with wavenumber $q=4\pi/5$ and amplitude $\varepsilon=0.1$ is added to give the sinusoidal profiles in (a)–(c) at $t=0$. Results in (d)–(f) show the solution of Equations (2a)–(2d) at $t=50$ where, in each case, we can see that each front moves in the positive $x$–direction and amplitude of the transverse perturbation clearly decays. Plots in (g)–(i) show the evolution of $\partial\Omega(t)$ at $t=0,2,4,6,8$ with the arrow showing the increasing value of $t$. Note that the plots in (g)–(i) focus on the interval $4\leq x\leq 6$ so that the details of the interface evolution are clear. Comparing these plots of $\partial\Omega(t)$ suggests that these travelling waves with $\kappa=-0.1$ are unstable as the perturbation amplitude appears to grow with time. All numerical calculations are performed on a $10\times 10$ domain uniformly discretised with a $201\times 201$ mesh.
@@ 3.4 Travelling wavefront stability: Linear stability analysis
were $q$ is the wave number and $\varepsilon\ll 1$ is the amplitude of the small perturbation Muller2002 ; Tam2023 . Here $\omega$ is an unknown constant that we will use to describe the short time decay $\omega<0$ or growth $\omega>0$ of the transverse perturbation amplitude. Our aim now is to assume that we have $\varepsilon\ll 1$ and use this small parameter to derive a perturbation solution to explore the dispersion relationship $\omega(q)$ for a range of travelling wave solutions characterised by different choices of $\kappa$ and $m$ which, as we showed in Section 3.2 implicitly defines the associated travelling wave speed $c$. This approach involves numerically solving a series of boundary value problems with one independent variable for the real component of the growth rate $\omega$.
@@
To compute these quantities we use linear interpolation to find positions $x_{j}$ such that $\phi(x,y_{j},t_{k})=0$ for the $j$th row in the discretisation of $\mathcal{D}$ at time $t=t_{k}$. The maximum and minimum values in the set ${x_{j}}$ determine $X_{\textrm{max}}(t)$ and $X_{\textrm{min}}(t)$, respectively. To estimate $\omega$ we use Equation (16) and assume that the perturbation amplitude grows or decays exponentially fast over a sufficiently short time interval. We use the Polynomials.jl package in Julia to perform a least–squares regression $\log(A)=\omega_{\textrm{n}}t+C$, where $C$ is a constant, and $\omega_{\textrm{n}}$ is the numerical estimate of the growth rate. For all results presented we consider a relatively short interval $t\in[1,2]$ with 21 equally–spaced time points within that interval to compute $\omega_{\textrm{n}}$. For each problem considered we generate a plot of $\log(A)$ as a function of $t$ which allows us to check that $A(t)$ grows or decays exponentially with time, as indicated by our data falling on a straight line over the time interval considered.
@@
Results in Figure 67 allow us to briefly return to Section 1 where, in Figure 1(e)–(h) where we hypothesized four potential outcomes in terms of invading, receding, linearly stable and linearly unstable travelling waves. Our results in Figure 78 indicate that invading travelling waves with $\kappa=0.1$ and $m=1$ are linearly stable while receding travelling waves with $\kappa=-0.1$ and $m=1$ are linearly unstable. We have found that these broad observations hold for other choices of $m$ and $\kappa$, namely that setting $\kappa>0$ leads to invading travelling waves that are linearly stable and setting $\kappa<0$ leads to receding travelling waves that are linearly unstable. Our open-source Julia codes provided on Github can be used to explore these other parameter choices. In summary, of the four potential outcomes previously hypothesized in Figure 1(e)–(h) we have only found evidence to support the scenario depicted in Figure 1(e) for $\kappa>0$ and Figure 1(h) for $\kappa<0$.
@@ 4 Conclusion and future work
In this work we re–visited the recently–established Porous–Fisher–Stefan (PFS) model of biological invasion, and focused our attention on properties of the travelling wave solutions of that model. The first presentation of the PFS model focused on working with a degenerate nonlinear diffusion term with flux $\mathcal{J}=-u\partial u/\partial x$ on $-\infty<x<L(t)$ with $\textrm{d}L(t)/\textrm{d}t=-\kappa u\partial u/\partial x$ at $x=L(t)$, and the existence of travelling waves describing biological invasion with $\kappa>0$ was established using phase plane and perturbation methods Fadai2020 . Here we extend this modelling framework by dealing with a 2D model that incorporates a generalised degenerate diffusion term with flux $-u^{m}\nabla u$, and we consider both invading fronts with $\kappa>0$, and receding fronts with $\kappa<0$. Solving the 2D model in an appropriate domain provides strong numerical evidence that the model supports long–time travelling wave solutions, and working in the phase plane allows us to examine properties of the travelling wave solutions over a range of model parameters in the PDE model, and to determine how the long–time travelling wave speed $c$ depends upon the choice of parameters in the PDE model.
@@
Preliminary explorations of the time–dependent PDE solutions indicate that invading travelling waves with $\kappa>0$ and $c>0$ are stable when a small transverse perturbation with wavenumber $q$ is introduced at the moving boundary, whereas receding travelling waves with $\kappa<0$ and $c<0$ are unstable. These informal numerical observations are corroborated with a linear stability analysis indicating that the short–time amplitude of the transverse perturbation increases when $\kappa<0$, whereas it decreases with $\kappa>0$. In general we find reasonably good agreement between the dispersion relationship obtained from the linear stability analysis and from the 2D level–set method, especially for small $q$. For larger $q$ it is straightforward to carry out the linear stability analysis because this involves solving two boundary value problems with just one independent variable that we discretise efficiently using a non-uniform mesh. In contrast, solving the 2D time–dependent PDE problem for larger $q$ requires an increasingly refined spatial mesh as $q$ increases, which makes the time–dependent PDE solutions impractical.
@@
As illustrated in Figures 78 setting $\kappa<0$ leads to linearly unstable travelling wave solutions. In the heat transfer literature, linearly unstable front propagation is often studied using regularisation by including additional mechanisms in the PDE model that can stabilise perturbation modes with large wavenumbers Chadam1983 ; Yang2002 . A standard approach in heat and mass transfer for Stefan–type moving boundary problems is to include terms describing surface tension effects Chadam1983 . Introducing surface tension involves replacing the usual 1D interface condition $u(L(t),t)=0$ with $u(L(t),t)=\gamma\mathcal{K}$, where $\gamma>0$ is the surface tension coefficient, and $\mathcal{K}$ is the signed curvature of the moving front Chadam1983 . This approach is reasonable in heat transfer and fluid mechanics where the dependent variable represents temperature and pressure, respectively. This means that the boundary condition $u(L(t),t)=0$ corresponds to a zero scaled temperature or zero scaled reference pressure, and it is physically reasonable to consider a negative scaled temperature or pressure at the 2D moving front with negative curvature. Working with the PFS model is very difference since here $u$ represents a scaled population density which means that it is physically unreasonable to set $u(L(t),t)=\gamma\mathcal{K}$ which would be negative at certain points on the 2D moving front with negative curvature. Therefore, in this work we have not pursued any regularisation of the linearly unstable travelling wave fronts as the classical approach is not suitable for this application.
@@
There are many opportunities for future work motivated by the problems that we have considered here. For example, one question for future consideration is to explore different options for biologically meaningful regularisation terms that could be applied both to linearly unstable front propagation for the PFS model, as well as other moving boundary problems that describe population biology phenomena. Other technical challenges could also be considered. For example, our phase plane analysis in Section 3.2 led to expressions for an exact solution for $c=0$ and $m=1$, which we used to construct perturbation solutions for $|c|\ll 1$. These solutions are insightful because they can be used to provide insight into the relationship between $\kappa$ and $c$ for $m=1$. Another option would be to use these solutions to construct perturbation solutions for $|m-1|\ll 1$ which could be used to provide insight into the relationship between $\kappa$ and $c$ for $m\neq 1$. We leave both of these extensions for future consideration.
@@
We solve Equations (19a)–(19d) by re–writing the boundary value problem in terms of $\phi_{0}=u_{0}^{(1+m)}$,
$\phi_{0}^{m/(1+m)}\left(\frac{\mathrm{d}^{2}\phi_{0}}{\mathrm{d}\xi^{2}}\right)+c\left(\frac{\mathrm{d}\phi_{0}}{\mathrm{d}\xi}\right)+(1+m)\phi_{0}\left(1-\phi_{0}^{1/(1+m)}\right)=0\;\;\text{on}\;\;-\infty<\xi<0,$ (23a)
$\lim_{\xi\to-\infty}\phi_{0}(\xi)=1,$ (23b)
$\phi_{0}=0\;\;\text{at}\;\;\xi=0,$ (23c)
$\frac{\mathrm{d}\phi_{0}}{\mathrm{d}\xi}$ $=-\frac{c(1+m)}{\kappa}\;\;\text{at}\;\;\xi=0.$ (23d)
@@
$\phi_{0}^{m/(1+m)}\left(\frac{\mathrm{d}^{2}\phi_{0}}{\mathrm{d}\xi^{2}}\right)+c\left(\frac{\mathrm{d}\phi_{0}}{\mathrm{d}\xi}\right)+(1+m)\phi_{0}\left(1-\phi_{0}^{1/(1+m)}\right)=0\;\;\text{on}\;\;-\xi_{\textrm{max}}<\xi<0,$ (24a)
$\phi_{0}=1\;\;\text{at}\;\;\xi=-\xi_{\textrm{max}}$ (24b)
$\phi_{0}=0\;\;\text{at}\;\;\xi=0,$ (24c)
$\frac{\mathrm{d}\phi_{0}}{\mathrm{d}\xi}$ $=-\frac{c(1+m)}{\kappa}\;\;\text{at}\;\;\xi=0,$ (24d)
@@
In general, $\phi_{0}$ varies rapidly with $\xi$ near $\xi=0$, whereas $\phi_{0}$ approaches unity as $\xi\to-\xi_{\textrm{max}}$. We take advantage of these properties by employing a non-uniform mesh where the mesh resolution is fine near $\xi=0$ and becomes increasingly coarse near $\xi=-\xi_{\textrm{max}}$. One way of achieving this is to use a geometrically spaced mesh McCue2022 where the mesh spacing varies from $\delta\xi_{\textrm{min}}$ at $\xi=0$ to $\delta\xi_{\textrm{max}}$ at $\xi=-\xi_{\textrm{max}}=-20$ using $N$ geometrically–spaced mesh points. At the $i$th internal mesh points we have $\delta\xi_{i}=(1+\epsilon)\delta\xi_{i+1}$. For $\delta\xi_{\textrm{min}}=1\times 10^{-5}$ with $N=301$ mesh points we obtain a geometric expansion factor of $(1+\epsilon)=1.0381894$ and $\delta\xi_{\textrm{max}}=0.7357$. Numerical experimentation confirms that this nonuniform mesh spacing leads to grid–independent results for the problems that we consider.
@@
At the $i$th internal mesh point we define $h^{+}_{i}=\xi_{i+1}-\xi_{i}$ and $h^{-}_{i}=\xi_{i}-\xi_{i-1}$ and discretise Equations(24a)–(24c) on the non-uniform mesh leading to a system of $N$ nonlinear algebraic equations $\boldsymbol{F}(\phi_{0}^{(i)})=\boldsymbol{0}$ which we solve using Newton-Raphson iteration. Elements of the vector $\boldsymbol{F}$ are given by,
$F_{0}$ $=\phi_{0}^{(0)}-1,$ (25a)
$F_{i}$ $={\phi_{0}^{(i)}}^{m/(1+m)}\left(A\phi_{0}^{(i+1)}+B\phi_{0}^{(i)}+C\phi_{0}^{(i-1)}\right)+c\frac{\phi_{0}^{(i+1)}-\phi_{0}^{(i-1)}}{h_{i}^{+}+h_{i}^{-}}+(1+m)\phi_{0}^{(i)}\left(1-{\phi_{0}^{(i)}}^{1/(1+m)}\right),$ (25b)
$F_{N}$ $=\phi_{0}^{(N)}\;,$ (25c)
@@
where the superscript $i=2,\dots,N-1$ refers to the $i$th internal mesh point, and the coefficients $A={2}/({h^{+}_{i}(h_{i}^{+}+h_{i}^{-})})$, $B={-2}/({h_{i}^{+}h_{i}^{-}})$, and $C={2}/({h^{-}_{i}(h_{i}^{+}+h_{i}^{-})})$ arise by applying standard finite difference stencils on the nonuniform mesh McCue2022 . An initial estimate of $\boldsymbol{\phi}_{0}^{(0)}$ can be updated using Newton-Raphson iteration as follows,
$\boldsymbol{\phi}_{0}^{(n+1)}=\boldsymbol{\phi}_{0}^{(n)}-J^{-1}\left(\boldsymbol{\phi}_{0}^{(n)}\right)\boldsymbol{F}\left(\boldsymbol{\phi}_{0}^{(n)}\right),$ (26)
@@
where the superscript $n$ gives the iteration count and $J_{ij}=\partial_{j}F_{i}$ are the elements of the Jacobian matrix for $\boldsymbol{F}(\phi_{0}^{(i)})$. Iterations are terminated when $\lVert\boldsymbol{\phi}_{0}^{(n+1)}-\boldsymbol{\phi}_{0}^{(n)}\rVert<1\times 10^{-6}$. In this work we take $\boldsymbol{\phi}_{0}^{(0)}$ to be a piecewise linear approximation of the shape of the travelling wave profile.
@@
To estimate the travelling wave speed $c$ we use a shooting method that commences with an initial estimate of $c^{(0)}$ to solve the leading order boundary value problem for $\phi_{0}(\xi)$. For our estimates of $c^{(0)}$ and $\phi_{0}(\xi)$ we discretise the moving boundary condition, Equation (23d), on our non-uniform mesh to give
$f^{(k)}=c^{(k)}+\frac{\kappa}{m+1}\left(P\phi_{0}^{({N-2},k)}-Q\phi_{0}^{({N-1},k)}+R\phi_{0}^{({N},k)}\right),$ (27)
@@
where $P=h^{+}_{N-1}/({h^{-}_{N-1}(h^{+}_{N-1}+h^{-}_{N-1})})$, $Q=(h^{+}_{N-1}+h^{-}_{N-1})/(h^{+}_{N-1}h^{-}_{N-1})$,, and $R=(2h^{+}_{N-1}+h^{-}_{N-1})/({h^{+}_{N-1}(h^{+}_{N-1}+h^{-}_{N-1})})$ is the estimate of the wave speed after $k$ iterations and while $\phi_{0}^{(i,k)}$ denotes our estimate of $\phi_{0}$ at the $i$th mesh point after $k$ iterations of the shooting method. To satisfy Equation (23d) we update our estimate of $c$ so that $f=0$, again using Newton-Raphson iteration,